Find the slope -intercept equation for the line with the indicated slope and y-intercept.
slope= -4
y-intercept= (0,-3/5)

Answer by mananth(16949)

m= -4

Plug value of the slope and point ( 0,- 3/5 ) in
Y = m x + b
-3/5 = -0 + b
b= -3/5 - -0
b= - 3/ 5
So the equation will be
Y = -4x - 3/5
m.ananth@hotmail.ca